OpenStudy (anonymous):

what is the solution of the equation 3/4X +5=-9

OpenStudy (haseeb96):

3+20x\4x =-9 3+20x =-36x 3+20x +36x=0 16x = -3 x=-3\16 this is the correct answer

OpenStudy (anonymous):

thanks

OpenStudy (haseeb96):

Wolud you not give me medal?